x-3y = 12
2y -3z = -9
2x +z = 7
solve for all three variables

Answer by jim_thompson5910(35256)

Hint:

Solve for z in 2x +z = 7 to get z = -2x+7

Then plug this into the second equation to get

2y -3z = -9

2y -3(-2x+7) = -9

2y + 6x - 21 = -9

2y + 6x = -9 + 21

2y + 6x = 12

6x + 2y = 12

-------------------------------------------------------

You now have the system

x-3y = 12
6x + 2y = 12

Use any method to solve this to find x and y. Then use x to find z.